# Improving the Plausibility of Pressure Distributions Synthesized from Depth through Generative Modeling

### v1

@@ 9 Model Parameters and Training Cost
10 Realtime Monitoring
To preliminarily evaluate real-time capability, we conducted a temporal analysis comparing two versions of our method: one trained on $8\,\mathrm{k}$ real samples and another on a larger $203\,\mathrm{k}$ mixed dataset. A qualitative visual comparison is presented in Fig. 7. A full video demonstration is available for reviewers only in the supplementary material.
We identified several limitations in the current real-time partial evaluation. First, while Fig. 7 demonstrates real-time capabilities, the current API (application-Interface) version is not optimized for speed. The current method does not explicitly enforce temporal consistency, which could lead to fluctuations between consecutive frames. A thorough zero-shot quantitative evaluation on unseen environments is needed. Addressing these aspects constitutes a clear direction for future work.
11 Additional Results
Figure 9 demonstrates the capability of ILS to generate diverse pressure distributions via anthropometric conditioning. Using a fixed depth input, the model synthesized three distinct pressure distributions: $\mathbf{\hat{p}}$ using actual parameters ($m=45.2,\ h=1.51,\ g=\text{Female}$), $\mathbf{\hat{p}_{2}}$ with reduced mass ($m=5.2,\ h=1.51,\ g=\text{Female}$), and $\mathbf{\hat{p}_{3}}$ with increased mass ($m=145.2,\ h=1.51,\ g=\text{Female}$). Reducing the input mass decreases the overall inferred pressure, while increasing it amplifies pressure, particularly in the pelvic region where high loads are expected. This adaptability suggests the system can accommodate patient weight fluctuations in real-world deployments by adjusting conditioning inputs, without other adjustments to the model.
The derived BM values are plotted in Fig. 9. While the BM calculated from $\mathbf{\hat{p}_{2}}$ and $\mathbf{\hat{p}_{3}}$ follows the trend of the conditioning inputs, it does not perfectly match the target values. This discrepancy arises because the input depth profile remains static, reflecting the subject’s original geometry—which conflicts with the synthetic mass prompts. Conversely, model depth - anthropometric relationship remain intact when using actual parameters $\mathbf{\hat{p}}$. Future work can lead to validating this method against ground-truth data from controlled weight gain and loss scenarios.
Qualitative results are shown in Fig. 10, including comparisons with LBBDM using a pretrained VQGAN. Complex scenarios, such as resting poses covered by a blanket, challenge all evaluated models. In these cases, baselines like BBDM and BPWnet often hallucinate incorrect postures. In contrast, the proposed method reduces error by blurring unseen regions rather than generating misleading features.
Figure 11 quantifies the MAE across supine, left-side, and right-side postures under three conditions. Consistent with expectations, the pelvic region exhibits the largest errors across all methods. Uncovered postures yield the lowest deviations, while scenarios involving a 3mm blanket result in the highest deviations due to occlusion. Proposed method and BBDM + ILS leads to lower MAE in pelvic region.
@@ 12 Model Configurations
LR Scheduler ReduceLROnPlateau
Scheduler Factor 0.5
Scheduler Patience 3000 (steps)
Min Learning Rate $5\times 10^{-7}$
@@
Conditional/ Unconditional discriminator PatchGAN
@@
Figure 10: Qualitative comparison of predicted pressure distributions across different methods. Rows represent distinct test samples, while columns display the outputs from competing models against the ground truth (GT). All models were trained on $8\,\mathrm{k}$ real samples.
@@
(b) training strategy of the denoising model in latent space
(c) schemetic diagram of the inference.
Figure 12: Schemetic of proposed LBBDM training strategy and inference.

### v2

@@ 10 Model Configurations
We remove the sigmoid activation from the final layer of AttnFnet+ILS in all experiments. Sigmoid produces bounded probability distributions that cause overestimation of both pressure values and body model BM parameters. Additionally, sigmoid saturation at extreme values compresses high/low pressure values into a narrow range, preventing the model from learning physically plausible pressure distributions.
@@
Cond/ Uncond discriminator PatchGAN
@@
11 Realtime Monitoring
To preliminarily evaluate real-time capability, we conducted a temporal analysis comparing two versions of our method: one trained on $8\,\mathrm{k}$ real samples and another on a larger $203\,\mathrm{k}$ mixed dataset. A qualitative visual comparison is presented in Fig. 8.
We identified several limitations in the current real-time partial evaluation. First, while Fig. 8 demonstrates real-time capabilities, the current API (Application Programming Interface) version is not optimized for speed. The current method does not explicitly enforce temporal consistency, which could lead to fluctuations between consecutive frames. A thorough zero-shot quantitative evaluation on unseen environments is needed. Addressing these aspects constitutes a clear direction for future work.
12 LBBDM Synthetic Pre-Training Results
This section presents results for a special case of LBBDM pre-training. Figure 7 illustrates the training strategy of the conditional LBBDM model as a schematic diagram. In contrast to the method described in Sec. 3.3, we tested the capability of conditional LBBDM to generate real samples when the AttnFnet model with ILS was pre-trained on synthetic data, while post-training and testing were performed exclusively on real samples.
Pre-training with synthetic data leads to an increase in MSE, as shown in Tab. 11. These results indicate the importance of real samples during pre-training. A visual comparison of both cases is presented in Fig. 11.
(b) training strategy of the denoising model in latent space
(c) schematic diagram of the inference.
Figure 7: schematic of proposed LBBDM training strategy and inference.
@@ 12 LBBDM Synthetic Pre-Training Results
Method Ucov cov (11mm) cov (33mm) overall
LBBDM $8\,\mathrm{k}$ 0.9047 1.0590 1.0803 1.0149
LBBDM-synth 1.0463 1.2318 1.2514 1.1765
Table 11: Test set MSE in $kPa^{2}$ for uncovered (Ucov), covered at 11 mm, covered at 33 mm, and overall poses. The results compare LBBDM pre-trained on real samples (LBBDM $8\,\mathrm{k}$) versus synthetic-only pre-training (LBBDM-synth).
13 Additional Results
Figure 10 demonstrates BM adaption capability of ILS via anthropometric conditioning. Using a fixed depth input, the model synthesized three distinct pressure distributions: $\mathbf{\hat{p}}$ using actual parameters ($m=45.2,\ h=1.51,\ g=\text{Female}$), $\mathbf{\hat{p}_{2}}$ with reduced mass ($m=5.2,\ h=1.51,\ g=\text{Female}$), and $\mathbf{\hat{p}_{3}}$ with increased mass ($m=145.2,\ h=1.51,\ g=\text{Female}$). Reducing the input mass decreases the overall inferred pressure, while increasing it amplifies pressure, particularly in the pelvic region where high loads are expected. This adaptability suggests the system can accommodate patient weight fluctuations in real-world deployments by adjusting conditioning inputs, without other adjustments to the model.
The derived BM values are plotted in Fig. 10. While the BM calculated from $\mathbf{\hat{p}_{2}}$ and $\mathbf{\hat{p}_{3}}$ follows the trend of the conditioning inputs, it does not perfectly match the target values. This discrepancy arises because the input depth profile remains static, reflecting the subject’s original profile—which conflicts with the synthetic mass prompts. Conversely, the model depth - anthropometric relationship remains intact when using actual parameters $\mathbf{\hat{p}}$. Future work can lead to validating this method against ground-truth data from controlled weight gain and loss scenarios.
Qualitative results are shown in Fig. 11, including comparisons with LBBDM with a pretrained VQGAN and LBBDM with pretraining on synthetic data. Complex scenarios, such as resting poses covered by a blanket, challenge all evaluated models. In these cases, baselines like BBDM and BPWnet often hallucinate incorrect postures. In contrast, the proposed method reduces error by blurring unseen regions rather than generating misleading features.
Figure 12 quantifies the MAE across supine, left-side, and right-side postures under three conditions. Consistent with expectations, the pelvic region exhibits the largest errors across all methods. Uncovered postures yield the lowest deviations, while scenarios involving a 3mm blanket result in the highest deviations due to occlusion. The proposed method and BBDM + ILS leads to lower MAE in the pelvic region.
@@ 13 Additional Results
Figure 11: Qualitative comparison of predicted pressure distributions across different methods. Rows represent distinct test samples, while columns display the outputs from competing models against the ground truth (GT). VQGAN represents pretraining happened with the VQGAN model on the CelebHQ dataset, synth represents AttnFnet-11M pretrained on synthetic data and post trainining preformed on real data.